The time taken to complete a particular type of a job is distributed approximately normally with mean 1.8 and standard deviation of 0. 1hours.Calculate the estimated completion time (to the nearest minute) that should be set so that there is a 90% chance that the job is completed on time                                                                                                 


1
Expert's answer
2021-10-14T14:24:51-0400

"P(Z<z)=0.9 \to z=1.28."

"\frac{x-1.8}{0.1}=1.28 \to x=1.8+1.28*0.1=1.928\; hrs=116\; min."